# LumiSculpt: A Consistency Lighting Control Network for Video Generation

### v1

@@ Abstract
Lighting plays a pivotal role in ensuring the naturalness of video generation, significantly influencing the aesthetic quality of the generated content. However, due to the deep coupling between lighting and the temporal features of videos, it remains challenging to disentangle and model independent and coherent lighting attributes, limiting the ability to control lighting in video generation. In this paper, inspired by the established controllable T2I models, we propose LumiSculpt, which, for the first time, enables precise and consistent lighting control in T2V generation models.LumiSculpt equips the video generation with strong interactive capabilities, allowing the input of custom lighting reference image sequences. Furthermore, the core learnable plug-and-play module of LumiSculpt facilitates remarkable control over lighting intensity, position, and trajectory in latent video diffusion models based on the advanced DiT backbone.Additionally, to effectively train LumiSculpt and address the issue of insufficient lighting data, we construct LumiHuman, a new lightweight and flexible dataset for portrait lighting of images and videos. Experimental results demonstrate that LumiSculpt achieves precise and high-quality lighting control in video generation.
@@
* Work done during internship at Ant Group. $\dagger$ Project lead.✉ Corresponding author.
Figure 1: LumiSculpt allows user-specified lighting intensity, direction, and trajectories, with textual conditions as inputs. Only one training session is needed for LumiSculpt to generate diverse results.
1 Introduction
“If a video tells a story, then lighting is the voice that shapes its tone and mood.”
@@ 1 Introduction
Lighting is essential for the essential of video generation, which is one of the defining factors for the overall aesthetic quality of the generated video, and is also used to convey emotions, highlight character traits, and guide the audience’s attention. Mainstream video generation methods currently employ latent diffusion models (LDMs) to achieve video generation through multi-step denoising in latent space. Research on controllable image and video generation based on LDMs supports our studies of consistent lighting control. Several methods (Ho et al., 2022b; Chen et al., 2023; Ho et al., 2022a; Wang et al., 2023a; Guo et al., 2024) have been developed to achieve relatively accurate text-controlled video generation, as well as video editing (Chai et al., 2023; Ceylan et al., 2023), customization (Wu et al., 2023), and controlling (Wang et al., 2023b; Wei et al., 2024; He et al., 2024). These works have improved the controllability, aesthetics, and usability of video generation. However, due to the deep coupling between lighting and the temporal features of videos, it is challenging to model independent and coherent lighting attributes, resulting in a lack of handy approaches to controlling lighting in videos.
@@
The challenge of customizing lighting lies in three aspects: the lack of training data, the representation of lighting, and the mechanism of injecting lighting features without influencing other attributes. Specifically, although there are currently relighting datasets based on light stages (Debevec et al., 2000), the data format of light stages is not easily applicable in video generation scenarios. Therefore, a flexible dataset that is adaptable to text-controlled new content generation is needed. Obtaining the projection of lighting on the camera’s imaging plane requires knowledge of the lighting and the surface texture of the illuminated object (Kim et al., 2024; Ren et al., 2024; Mei et al., 2024), which cannot be satisfied in an end-to-end video generation scenario. Thus, a simpler lighting representation that is only related to lighting parameters is important. Finally, similar to most control tasks, lighting control faces the problem of the deep decoupling of lighting from other elements, such as semantics and color.
@@
In this paper, we propose LumiHuman to solve the problem of limited training data. LumiHuman is a portrait lighting dataset that can constitut more than 220K videos of humans with known lighting parameters. This is a lightweight and flexible dataset that is not limited to specific lighting movements but is presented in freely combinable frames, laying the foundation for a more diverse range of lighting paths and combinations. We then use virtual engine rendering with known lighting parameters to obtain projections of different directional lighting on planes as a lighting representation. Additionally, we propose a consistency lighting control network, LumiSculpt, which learns an accurate plug-and-play lighting module capable of controlling the direction and movement of lighting in video generation. To solve the problem of lighting feature injection, we introduce a light control module that takes the lighting projection as input and integrates lighting control injected into the generative model layer by layer. Furthermore, to better decouple lighting and appearance, we design a decoupling loss based on a dual-branch structure, preserving diverse generative capabilities. We implemented LumiSculpt on Open-Sora (Lab & etc., 2024) to enable precise lighting control. We conducted comprehensive quantitative and qualitative evaluations. The experimental results show that LumiSculpt has achieved state-of-the-art performance in the control of text-to-video lighting, as shown in Figure 1. In summary, our main contributions are as follows:
@@
  • We propose a portrait lighting dataset LumiHuman, which includes 65 different individuals and allows for the free combination of over 3000 lighting trajectories. LumiHuman paves the way for more lighting control in both image and video generation.
@@
  • We introduce LumiSculpt, enabling control of lighting direction and movement trajectories in video generation. We propose a lighting representation method, a lighting injection approach, and a lighting decoupling loss for the text-to-video generation scenario, enabling diverse content generation with limited data.
  • Extensive experiments prove that LumiSculpt has achieved state-of-the-art performance. LumiSculpt can not only generate accurate lighting intensity, direction, and trajectories but also maintain coherence and content diversity.
2 Related Works
2.1 Relighting
@@
In recent years, deep learning techniques have made significant progress in portrait relighting (Kim et al., 2024; Mei et al., 2023; 2024; Nestmeyer et al., 2020; Pandey et al., 2021; Sun et al., 2019; Wang et al., 2020; Yeh et al., 2022; Zhang et al., 2021), often relying on paired data captured by light stage systems (Debevec et al., 2000) for supervised learning. Typically, these methods require the use of high dynamic range (HDR) environmental maps as input. This process involves estimating intermediate surface properties, including normal vectors, albedo, diffuse reflectance, and specular reflection characteristics. However, the reliance on HDR environmental maps limits the practical application of these techniques in video generation scenarios. Besides, researchers also explore portrait relighting techniques that do not depend on light stage data (Hou et al., 2021; 2022; Wang et al., 2023c).
@@
Recently, diffusion-based models have brought new research directions to the field. Ren et al. (Ren et al., 2024) propose a three-stage lighting-aware diffusion model called Relightful Harmonization, which aims to provide complex lighting coordination for foreground portraits with any background image. Zeng et al. (Zeng et al., 2024) propose a three-stage portrait relighting method using a fine diffusion model called DiLightNet, which calculates radiance cues to re-synthesize and refine the foreground object by combining the rough shape of the foreground object inferred from the preliminary image. Xing et al. (Xing et al., 2024) propose a natural image relighting method called Retinex-Diffusion, which treats the diffusion model as a black-box image renderer and strategically decomposes its energy function to be consistent with the image formation model. However, there is still a lack of methods for lighting control in text-to-video generation.
2.2 Text-to-video synthesis and controlling
@@
Recently, several researches, such as (Ho et al., 2022b; Chen et al., 2023; Ho et al., 2022a; Wang et al., 2023a; Guo et al., 2024), have adopted diffusion models to create highly realistic video content, utilizing text as conditions in guiding the generation process. These studies focus on ensuring consistency between textual descriptions and the final video output. Addressing the issue of difficulty in precisely describing specific visual attributes through text conditions, some studies have attempted to achieve finer video control by fine-tuning models or introducing additional control parameters. Tune-A-Video (Wu et al., 2023) propose a fine-tuning framework that allows users to customize specific videos. VideoComposer (Wang et al., 2023b) use explicit control signals to guide the temporal dynamics of the video. Gong et al. (Gong et al., 2023) introduce TaleCrafter to handle interactions among multiple characters, featuring layout and structural editing capabilities. He et al. (He et al., 2023) propose a retrieval-based deep guidance method that can integrate existing video clips into a coherent narrative video by customizing the appearance of characters. These studies mainly focus on the appearance of visual content or objects. Several methods (Zhao et al., 2023; Zhang et al., 2023; Wei et al., 2024; He et al., 2024) learn and controll motion through customized diffusion models. These attempts have made pioneering progress in controlling video in specific aspects. However, there is still a lack of effective solutions for precisely controlling the lighting effects in videos.
3 Methods
@@
Our goal is to achieve unified control over video lighting, a challenging task with considerable implications. The primary difficulties are threefold: (1) Dataset Scarcity: there is a significant lack of lighting datasets, particularly for videos, with few annotated examples that explicitly showcase lighting changes and where lighting information is well-defined. (2) Complexity of Lighting Attributes: lighting encompasses various factors, including the type of light source, direction of illumination, and the material properties of objects. In the context of text-to-video generation, the material and shape of the generated objects are often unknown. Thus, effectively representing lighting information to accurately convey its visual effects within the camera’s field of view poses a substantial challenge. (3) Attribute Decoupling: like many control tasks, lighting control faces the challenge of decoupling specific attributes. A key technical hurdle is how to isolate lighting information from the appearance of the training data, ensuring that the model does not overfit to the training data’s appearance or incorporate irrelevant details.
Thus, this section is organized as follows. Section 3.1 presents the preliminary content of the video generation models; Section 3.2 introduces the method of collecting the lighting dataset that we propose, namely LumiHuman. Section 3.3 introduces the method of representing camera information and injecting it into the text-to-video model, namely LumiSculpt.
3.1 Preliminary
Text-to-video diffusion models
Recently, significant advancements have been made in Text-to-Video (T2V) diffusion models (Chen et al., 2023; Wang et al., 2023a; Lab & etc., 2024; Hong et al., 2022). Most of these models follow the classic algorithmic framework from the field of image generation. Specifically, the process typically involves gradually introducing noise $\epsilon$ into $N$ sequences of images $z_{1},\ldots,z_{N}$ until they approximate a Gaussian distribution. Given noise-corrupted inputs $z_{1},\ldots,z_{N}$, a neural network is trained to predict the added noise. During training, the network strives to minimize the mean squared error (MSE) between its noise predictions and the actual noise.
Figure 2: The collection process of LumiHuman includes: (a) designing a 3D point light source matrix, (b) rendering single-frame images and generating portrait lighting videos with various path lighting and lighting reference videos, (c) annotating with a BLIP model, and (d) producing enhanced background captions using a large language model.
3.2 LumiHuman
Lighting Videos
Extracting lighting information from existing videos presents significant challenges, even for professionals. The complexity of lighting encompasses factors such as light source type, angle, and distance. Consequently, annotating existing video data to create lighting datasets is challenging. Generating data with known lighting parameters through generative methods is a more viable approach.
We introduce a portrait lighting dataset, referred to as LumiHuman. This dataset is created using a unreal engine for lighting simulation, allowing for the production of data with known lighting information. The LumiHuman collection comprises five key stages: (1) Lighting Design: As illustrated in Figure 2(a), we developed a lighting position matrix, i.e., a three-dimensional grid measuring $160cm\times 160cm\times 160cm$. The points are uniformly spaced at $5cm$ intervals to serve as lighting positions. Point light sources move within these grid points to capture data of the subject being illuminated from various angles. (2) Lighting Trajectories Design: Within the three-dimensional grid, we defined horizontal, vertical, and diagonal trajectories, each composed of grid points to simulate diverse lighting change effects. (3) Character Construction: To facilitate the production of portrait lighting data, we utilized the MetaHuman dataset (MetaHuman, 2023), which features 3D models of 65 different individuals. This diversity enhances the visual effects of light projection across different characters. (4) Flexibility and Storage Optimization: To achieve a wide variety of lighting path variations while addressing storage concerns, i.e., given the presence of duplicate frames at the same lighting positions, we offer a flexible and lightweight image-video dataset. This dataset includes images rendered from various lighting positions in a three-dimensional grid, where each character corresponds to $33\times 33\times 3$ 33 different sampled images. In practice, videos can be generated from images in the dataset using predefined paths, or additional paths can be designed to simulate different lighting effects, as shown in Figure 2 (b). (5) Text Annotation and Augmentation: For automatic text annotation of the videos, we employed BLIP (Li et al., 2022). Additionally, we utilized GPT-4 (Achiam et al., 2023) for caption augmentation, generating diverse contextual backgrounds for the dataset, as shown in Figures 2(c) and (d).
@@ 3.2 LumiHuman
To describe the effect of the lighting projection, a straightforward approach is to incorporate the lighting parameters as additional information in the model. However, this method requires a substantial amount of annotated data to establish a mapping between the lighting vectors and the two-dimensional plane. To better align with the model’s inferential feature space, we propose projecting lighting information into an empty space, as illustrated in Figure 3(b). For different lighting positions, this is represented as an image where brighter areas indicate stronger illumination, and darker areas signify weaker lighting. This representation allows for a more effective alignment of lighting information with the video generation model.
@@
Figure 3: The pipeline of LumiSculpt consists of the generation backbone, i.e. the controlled branch, which includes (a) a diffusion transformer (DiT), a pre-trained video denoising network, and (b) a light encoder, a trainable external transformer network. The light encoder takes light reference latents as input and processes them through various blocks to produce a light condition sequence. This sequence is integrated into the generation backbone using several (d) merge modules within each block. During training, aside from the controlled branch, (c) a frozen branch is introduced to provide regularization for diverse appearances. The frozen branch is a DiT with frozen parameters, sharing weights with (a). Both branches predict noise, resulting in $\epsilon_{pred}$ and $\epsilon_{reg}$, which are used to compute the disentanglement loss $\mathcal{L}_{dis}$.
3.3 LumiScuplt
3.3.1 Integrating Lighting into Video Generators
Since lighting is represented in pixel space, it can be parameterized as input to the standard visual model. We extract lighting features using a Variational Autoencoder (VAE) shared with a generative model, which are then fed into a dedicated lighting encoder. As illustrated in Figure 3(b), this lighting encoder employs a transformer architecture comprised of self-attention layers, enabling it to compute attention scores globally across the video. This design allows the encoder to effectively capture the spatial and temporal relationships of lighting throughout the video clip. The lighting encoder accepts lighting features as input. The transformer blocks, matching the number of layers in the backbone model, output a sequence of latents of the same size. For each layer in the backbone model, the lighting encoder provides corresponding features of identical dimensions, facilitating feature fusion. Our objective is to seamlessly integrate these latents into the DiT architecture of the T2V model, as shown in Figure 3(d). The latent features of the video, denoted as $z_{t}$, and the lighting features, $c_{t}$, are combined through element-wise addition. This integrated feature is then passed through a linear layer, producing the output for the subsequent layer with a hyper-parameter termed guidance scale set to 0.5.
3.3.2 Lighting Learning
We utilize a data-driven approach to learn complex lighting information, capturing the projection effects of lighting at different positions on the human face. The implementation of this method is based on the LumiHuman. There is an issue with simply learning lighting from data rendered by unreal engines, which is the leakage of appearance information. This occurs due to the consistent backgrounds and layouts in the dataset. While such consistency provides more stable training data for lighting learning, it simultaneously increases the model’s susceptibility to overfitting to similar appearances. This raises the key issue of lighting control: how to decouple lighting from other elements? We propose a lighting-appearance disentanglement method, which includes a dual-branch framework and a novel disentanglement loss.

### v2

# LumiSculpt: Enabling Consistent Portrait Lighting in Video Generation

@@ Abstract.
Lighting plays a pivotal role in ensuring the naturalness and aesthetic quality of video generation. However, the impact of lighting is deeply coupled with other factors of videos, e.g., objects and scenes. Thus, it remains challenging to disentangle and model coherent lighting conditions independently, limiting the flexibility to control lighting in video generation. In this paper, inspired by the established controllable T2I models, we propose LumiSculpt , which enables precise and consistent lighting control in T2V generation models. LumiSculpt equips the video generation with new interactive capabilities, allowing the input of reference image sequences with customized lighting conditions. Furthermore, the core learnable plug-and-play module of LumiSculpt facilitates direct control over the intensity, position and trajectory of an assumed light source in video diffusion models. To effectively train LumiSculpt and address the issue of insufficient lighting data, we construct LumiHuman , a new lightweight and flexible dataset for portrait lighting of images and videos. Experimental results demonstrate that LumiSculpt achieves precise and high-quality lighting control in video generation. The analysis demonstrates the flexibility of LumiHuman .
Light Control, Video generation, Text-to-video generation, Diffusion models
$c$s: Computing methodologies Computer vision
Figure 1. LumiSculpt allows user-specified control over the intensity, position, and trajectories of an assumed light source, with textual conditions as input. Being trained once, LumiSculpt is capable of generating diverse results at inference time.
@@
* Work done during internship at Ant Group. $\dagger$ Project lead. Corresponding author.
1. Introduction
@@
If a video tells a story, then lighting is the voice that shapes its tone and mood. Lighting is essential for video generation, which is one of the defining factors for the overall aesthetic quality of the generated video, and is also used to convey emotions, highlight character traits, and guide the audience’s attention. Mainstream video generation methods currently employ latent diffusion models (LDMs) to achieve video generation through multi-step denoising in a latent space. Research on controllable image and video generation based on LDMs supports our studies of consistent lighting control. Several methods (Ho et al., 2022b; Chen et al., 2023; Ho et al., 2022a; Wang et al., 2023b; Guo et al., 2024) have been developed to achieve relatively accurate text-controlled video generation, as well as video editing (Chai et al., 2023; Ceylan et al., 2023), customization (Wu et al., 2023), and controlling (Wang et al., 2023c; Wei et al., 2024; He et al., 2024). These works have improved the controllability, aesthetics, and usability of video generation. However, due to the deep coupling between lighting and the other non-stationary factors of videos, it is challenging to model coherent lighting conditions independently, resulting in a lack of handy approaches to controlling lighting in videos.
@@
The challenge of customizing lighting lies in three aspects: the lack of training data, the representation of lighting, and the mechanism of injecting lighting conditions without influencing other attributes. Specifically, although there are relighting datasets based on light stages (Debevec et al., 2000), the data format of light stages is not readily applicable in video generation scenarios. Therefore, a new dataset that is adaptable to text-controlled content generation is needed. Obtaining the projection of lighting on the camera’s imaging plane requires knowledge of the lighting and the surface texture of the illuminated objects (Kim et al., 2024; Ren et al., 2024; Mei et al., 2024), which cannot be achieved in an end-to-end video generation scenario. Thus, a simple lighting representation that is only related to lighting parameters is important. Finally, similar to most control tasks, lighting control faces the problem of the deep decoupling of lighting from other factors, such as semantics and color.
@@
In this paper, we propose LumiHuman to address the problem of limited training data. LumiHuman is a portrait lighting dataset that can constitut more than 220K videos of humans with known lighting parameters. This is a lightweight and flexible dataset that is not limited to specific lighting movements but is presented in freely combinable frames, laying the foundation for a more diverse range of lighting paths and combinations. We then use virtual engine rendering with known lighting parameters to obtain projections of different directional lighting on planes as a lighting representation. To achieve video lighting control, we propose a lighting control method, LumiSculpt , which learns an accurate plug-and-play lighting module capable of controlling the direction and movement of lighting in video generation. To solve the problem of lighting feature injection, we introduce a light control module that takes the lighting projection as input and integrates lighting control injected into the generative model layer by layer. Furthermore, to better decouple lighting from appearance, we design a decoupling loss based on a dual-branch structure, preserving diverse generative capabilities.
We implemented LumiSculpt  on Open-Sora-Plan (Lab and etc., 2024) to enable precise lighting control. LumiSculpt focuses specifically on lighting control in text-to-video generation, rather than addressing image-to-image or video-to-video relighting tasks. It goes beyond managing fixed types and directions of lighting based on text prompts. Instead, it enables precise control over the position and moving trajectory of the light source within a single video, offering great flexibility and customization. We conducted comprehensive quantitative and qualitative evaluations. The experimental results show that LumiSculpt  has achieved state-of-the-art performance in the control of text-to-video lighting in precision and diversity, as shown in Figure 1. In summary, our main contributions are as follows:
@@
  • We introduce a portrait lighting dataset LumiHuman . LumiHuman is a lighting video dataset comprising over 220K different videos (i.e., 2.3M images). LumiHuman includes over 30K lighting positions, and over 3K light source trajectories for each individual. LumiHuman paves the way for lighting control in both image and video generation.
@@
  • We introduce LumiSculpt , which enables control of the position and moving trajectories of light sources in video generation. We propose a lighting representation method, a lighting injection approach, and a lighting decoupling loss for the text-to-video generation scenario, which enables diverse content generation with limited data.
  • Extensive experiments prove that LumiSculpt has achieved state-of-the-art performance. LumiSculpt can enable consistent portrait lighting in diverse scenarios. The analysis demonstrate the comprehensiveness and flexibility of LumiHuman , which provides a generalizable lighting prior.
2. Related Works
2.1. Relighting
@@
In recent years, deep learning techniques have made significant progress in portrait relighting (Kim et al., 2024; Mei et al., 2023, 2024; Nestmeyer et al., 2020; Pandey et al., 2021; Sun et al., 2019; Wang et al., 2020; Yeh et al., 2022; Zhang et al., 2021), which often rely on paired data captured by light stage systems (Debevec et al., 2000) for supervised learning. Typically, these methods require the use of high dynamic range (HDR) environmental maps as input. This process involves estimating intermediate surface properties, including normal vectors, albedo, diffuse reflectance, and specular reflection characteristics. However, the reliance on HDR environmental maps limits the practical application of these techniques in video generation scenarios. Besides, researchers also explore portrait relighting techniques that do not depend on light stage data (Hou et al., 2021, 2022; Wang et al., 2023a).
@@
Recently, diffusion-based models have shed light on new approaches to relighting. Ren et al. (Ren et al., 2024) propose a three-stage lighting-aware diffusion model called Relightful Harmonization, which aims to provide complex lighting coordination for foreground portraits with any background image. Zeng et al. (Zeng et al., 2024) propose a three-stage portrait relighting method using a fine diffusion model called DiLightNet, which calculates radiance cues to re-synthesize and refine the foreground object by combining the rough shape of the foreground object inferred from the preliminary image. Xing et al. (Xing et al., 2024) propose a natural image relighting method called Retinex-Diffusion, which treats the diffusion model as a black-box image renderer and decomposes its energy function to be consistent with the image formation model. However, there still lack of methods for lighting control in text-to-video generation. The most related works are diffusion-based relighting methods. LightIt (Kocsis et al., 2024) is an image-guided method for image relighting conditioning on shading estimation and normal maps. IC-Light (Zhang et al., 2024) is an image relighting method to generate harmonized background with the user input foreground. Light-A-Video (Zhou et al., 2025) proposes a training-free video relighting method, which leverages on the image relighting model to achieve video relighting controlled by text. It is worth noting that the above methods are image-to-image or video-to-video methods that can leverage the input prior, while our method aims at lighting control in text-to-video generation, which is more challenging. Moreover, our method is able to control the exact position and trajectory of the light source within a single video, instead of setting a fixed light type using text prompts.
Figure 2. Diverse indivisuals in LumiHuman .
Table 1. Comparison of Openillumination (Liu et al., 2024),DPR (Zhou et al., 2019) and LumiHuman
Dataset Synthesis Light Positions Light Movement Number of Images Subject Resolutions
DPR 2D 7 None 138K - 1024×1024
Openillumination Light Stage 142 None 108K 64 objects 3000×4096
LumiHuman 3D 35,937 ¿3K 2.3M 65 individuals 1024×1024
2.2. Text-to-video synthesis and controlling
@@
Recently, several researches, such as (Ho et al., 2022b; Chen et al., 2023; Ho et al., 2022a; Wang et al., 2023b; Guo et al., 2024), have adopted diffusion models to create highly realistic video content, utilizing text as conditions in guiding the generation process. These studies focus on ensuring alignment between textual descriptions and the final video output. Addressing the issue of difficulty in precisely describing specific visual attributes through text conditions, some studies have attempted to achieve finer video control by fine-tuning models or introducing additional control parameters. Tune-A-Video (Wu et al., 2023) propose a fine-tuning framework that allows users to customize specific videos. VideoComposer (Wang et al., 2023c) use explicit control signals to guide the temporal dynamics of the video. Gong et al. (Gong et al., 2023) introduce TaleCrafter to handle interactions among multiple characters, featuring layout and structural editing capabilities. He et al. (He et al., 2023) propose a retrieval-based deep guidance method that can integrate existing video clips into a coherent narrative video by customizing the appearance of characters. These studies mainly focus on the appearance of objects and scenes. Several methods (Zhao et al., 2023; Zhang et al., 2023b; Wei et al., 2024; He et al., 2024) learn and control motion through customized diffusion models. These attempts have made substantial progress in controlling video in specific aspects. Nevertheless, further dedicated efforts are required on precise control of lighting in videos.
3. LumiHuman
@@
Our goal is to achieve unified control over video lighting—a challenging task with significant implications. The primary difficulties can be categorized into three areas: (1) Dataset Scarcity: There is a notable lack of lighting-specific datasets, particularly for videos. Few annotated examples explicitly capture lighting variations, and even fewer provide well-defined lighting information. (2) Complexity of Lighting Attributes: Lighting involves multiple factors, including light source type, direction of illumination, and the material properties of objects. Accurately representing lighting effects within the camera’s field of view becomes especially important. (3) Attribute Decoupling: Similar to other control tasks, lighting control requires effective decoupling of specific attributes. A major technical challenge lies in isolating lighting information from object appearance in training data, preventing the model from overfitting.
We introduce a portrait lighting dataset, referred to as LumiHuman . LumiHuman is a continuous lighting video dataset comprising over 220K different videos (i.e., 2.3 million images). The resolution of each video is $1024\times 1024$. LumiHuman is created using Unreal Engine (Epic Games, 2024) for lighting simulation, allowing for the production of data with known lighting information. As shown in Figure 2, LumiHuman includes 65 diverse human subjects, 30K lighting positions, and over 3K lighting trajectories for each people. As shown in Table. 1, compared to other lighting datasets Openillumination (Liu et al., 2024) and Deep Portrait Relighting (DPR) dataset (Zhou et al., 2019) (generated from face image dataset Celeb-A (Liu et al., 2015)), LumiHuman outperforms in light positions, light movements and number of images.
Figure 3. (a) LumiHuman offers a variety of basic elements that can be combined to form various types of portrait lighting, widely applicable to a range of tasks related to character lighting. (b) shows the distribution of light intensity on different facial areas of the characters; LumiHuman ’s lighting matrix can cover all areas of the face and produce a significant range of light and shadow variations. (c) shows an example of creating a continuous lighting video using LumiHuman .
Details
As shown in Figure 3(a), LumiHuman can be combined to generate various types of character lighting. Its 30K lighting positions enable the creation of light and shadow effects across all areas of the human face. In Figure 3(b), we present the brightness distribution map for different facial regions. Each ridge in the ridge plot represents a specific facial area, where the horizontal axis indicates brightness and the vertical axis denotes the number of samples corresponding to each brightness level. LumiHuman comprehensively covers all facial areas and distributes samples across a wide range of brightness levels. As illustrated in Figure 3(c), we present continuous video frames composed of samples from LumiHuman . These samples can be flexibly arranged to form diverse lighting trajectories based on user specifications—such as horizontal, vertical, diagonal, arc-shaped, or multi-light-source superpositions.
@@
A straightforward approach to representing lighting effects is to embed lighting parameters as additional input to the model. However, this method requires a large amount of annotated data to establish a reliable mapping between lighting vectors and the two-dimensional image plane. To better align with the model’s latent space, we propose projecting lighting information into a blank canvas, as illustrated in Figure 5(b). For each lighting position, this is visualized as an image in which brighter regions indicate stronger illumination, and darker regions indicate weaker lighting. This representation enables more effective integration of lighting information into the video generation model.
Figure 4. The collection process of LumiHuman includes: (a) designing a 3D point light source matrix of $33\times 33\times 33$ lighting points, (b) rendering single-frame images and generating portrait lighting videos with various path lighting and lighting reference videos, (c) annotating with a BLIP model, and (d) producing enhanced background captions using a large language model.
Data Collection
The LumiHuman  collection consists of five key stages: (1) Lighting Design: As illustrated in Figure 4(a), we constructed a lighting position matrix—a three-dimensional grid measuring $160\,\text{cm}\times 160\,\text{cm}\times 160\,\text{cm}$. Points within the grid are uniformly spaced at $5\,\text{cm}$ intervals and serve as lighting positions. A point light source moves across these grid points to illuminate the subject from diverse angles. (2) Lighting Trajectory Design: Within this 3D grid, we defined horizontal, vertical, and diagonal trajectories composed of grid points to simulate a variety of lighting change effects. (3) Character Construction: To generate high-quality portrait lighting data, we employed the MetaHuman dataset (MetaHuman, 2023), which provides 3D models of 65 diverse individuals. This variation enhances the richness of visual lighting effects across different characters. (4) Flexibility and Storage Optimization: To support a wide range of lighting path variations while managing storage demands—particularly due to duplicate frames from identical lighting positions—we provide a flexible, lightweight image-video dataset. The dataset includes images rendered from different lighting positions within the 3D grid, with each character associated with $33\times 33\times 33$ unique samples. Videos can be generated using predefined lighting trajectories, or users can define new paths to simulate additional lighting effects, as demonstrated in Figure 4(b). (5) Text Annotation and Augmentation: For automated video captioning, we utilized BLIP (Li et al., 2022). To enrich the contextual diversity of the dataset, we further applied GPT-4 (Achiam et al., 2023) for caption augmentation, generating a wide range of descriptive narratives, as illustrated in Figures 4(c) and (d).
@@
Figure 5. The pipeline of LumiSculpt consists of the generation backbone, i.e. the controlled branch, which includes (a) a diffusion transformer (DiT), a pre-trained video denoising network, and (b) a light encoder, a trainable external transformer network. The light encoder takes light reference latents as input and processes them through various blocks to produce a light condition sequence. This sequence is integrated into the generation backbone using several merge modules within each block. During training, we propose a (c) dual-branch framework including a controlled branch and a frozen branch, which provide regularization for diverse appearances. The frozen branch is a DiT with frozen parameters, sharing weights with (a). Both branches predict noise, resulting in $\epsilon_{pred}$ and $\epsilon_{reg}$, which are used to compute the disentanglement loss $\mathcal{L}_{dis}$. (c) and (d) show that LumiSculpt differs form ControlNet (Zhang et al., 2023a) in terms of model structure, condition injection, training manners and objectives.
4. LumiSculpt
4.1. Integrating Lighting into Video Generators
LumiHuman enables lighting to be represented in pixel space and parameterized as an input to standard visual models. We extract lighting features using a Variational Autoencoder (VAE), shared with the generative model, and feed them into a dedicated lighting encoder. As illustrated in Figure 5(b), the lighting encoder adopts a transformer architecture composed of self-attention layers, enabling it to compute global attention scores across video frames. This design effectively captures the spatial and temporal dynamics of lighting throughout a video clip.
The lighting encoder takes the lighting features as input and passes them through transformer blocks that match the number of layers in the backbone model. These blocks output a sequence of latent representations with dimensions aligned to those of the backbone model, enabling seamless feature fusion. Our objective is to integrate these latents into the DiT architecture of the T2V model. Specifically, the latent video features, denoted $z_{t}$, are combined with the lighting features $c_{t}$ via element-wise addition. The resulting features are then passed through a linear layer to produce the output for the next layer, with a guidance scale hyperparameter set to 0.5.
As shown in Figure 5(d), LumiSculpt employs a 3D self-attention mechanism in the lighting encoder and utilizes multi-stage weighting as the conditional injection mechanism. In comparison, ControlNet uses a U-Net encoder to extract features and injects conditions via additive latent fusion.
4.2. Lighting Learning
We adopt a data-driven approach to learn complex lighting effects, specifically capturing the impact of light projection from various positions on the human face. This approach is implemented using LumiHuman . However, a key challenge arises when learning lighting directly from data rendered in Unreal Engine: the leakage of appearance information. This issue stems from the dataset’s consistent backgrounds and layouts, which, while beneficial for stable training, also increase the risk of the model overfitting to specific appearances.
This highlights a central challenge in lighting control: how can lighting be disentangled from other visual attributes? We address this by proposing a lighting-appearance disentanglement method, which incorporates a dual-branch architecture alongside a novel disentanglement loss function. This framework enables the model to isolate lighting information while mitigating the influence of appearance-related features.